OpenStudy (bethb):

how do I draw this A quadrilateral has vertices A(11, -7), B(9, -4), C(11, -1), and D(13, -4). Quadrilateral ABCD is a...... If the vertex C(11, -1) were shifted to the point C′(11, 1), quadrilateral ABC′D would be a.....

OpenStudy (bethb):

part 1. A. is a parallelogram with nonperpendicular and noncongruent adjacent sides B. trapezoid with exactly one pair of parallel sides C. rectangle with noncongruent adjacent sides D. rhombus with nonperpendicular adjacent sides . part 2 A. parallelogram with nonperpendicular adjacent sides B. kite C. square D. rhombus with nonperpendicular adjacent sides .

OpenStudy (eliesaab):

Part1 Compute the distances AB, BC, CD, CA If they are equal what Can you conclude?

OpenStudy (eliesaab):

Part1 Compute the slopes of AB, BC,CD, DA and see what you can get.

OpenStudy (bethb):

if they are equal then its a square?

OpenStudy (eliesaab):

\[ AB=\sqrt{(-4+7)^2+(9-11)^2}=\sqrt{13} \] Do the same for the rest

OpenStudy (eliesaab):

No, we can only conclude that it is rhombus. If in addition you can show that one of the angles is 90 degrees, then it would be a circle

OpenStudy (eliesaab):

I do the slope of AB and you do the rest. \[ s_{AB}=\frac{-7+4} {11-9}=-\frac32 \]

OpenStudy (eliesaab):

Do them and let us discuss what you can conclude

OpenStudy (bethb):

I got AC= 8 BC= 3 CD=3 AD= sqrt13

OpenStudy (eliesaab):

All of the sides are equal to \[ \sqrt{13} \]

OpenStudy (bethb):

im doing the slope now

OpenStudy (eliesaab):

See the graph

OpenStudy (bethb):

wait when you add them they equal sqrt13 or all of them are sqrt13

OpenStudy (eliesaab):

Every side is equal \( \sqrt{13}\)